rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#9969 — FS#13824 — Seattle

Attached to Project— Network
Incident
Backbone US
CLOSED
100%
Our router in unreachable. We will investigate.

The Seattle CDN node and Anycast DNS are impacted.
Date:  Thursday, 02 July 2015, 12:19PM
Reason for closing:  Done
Comment by OVH - Wednesday, 01 July 2015, 13:48PM

Fiber Cut

Level 3 Communications

This is to confirm that an update has been recorded on Level 3 ticket 9409299.

Summary
A suspected fiber cut is affecting unprotected services in Portland , OR.


Event Case ID: 9408548
Location: Portland, OR
Impacted For: 1 hour 27 minutes
ETR: Unknown
Bridge: Blue
Updates

16:13 GMT - The fiber cut has been located and was found to be caused by a contractor drilling. A splice crew has been engaged to dispatch and is expected to arrive onsite at 17:00 GMT.

15:56 GMT - Field technicians are walking the route to locate the physical location of the damage. No additional information is available at this time.

15:31 GMT - The Transport NOC advised that a suspected fiber cut is affecting unprotected services in Portland, OR. A field technician was engaged and was able to shoot the fiber 1100ft from the gateway. The field technician is now looking for the exact location of the fault at this time.


Comment by OVH - Wednesday, 01 July 2015, 13:48PM

*** CASCADED EXTERNAL NOTES 30-Jun-2015 17:44:30 GMT From CASE: 9408548 - Event
The Transport NOC reported that crews are proofing the conduit at this time to ensure that new cable can be pulled though. Restoration efforts continue.


Comment by OVH - Wednesday, 01 July 2015, 13:48PM

*** CASCADED EXTERNAL NOTES 30-Jun-2015 19:28:30 GMT From CASE: 9408548 - Event
The spare cable arrival has been delayed. A new arrival time of 20:30 GMT was provided.


Comment by OVH - Wednesday, 01 July 2015, 13:48PM

*** CASCADED EXTERNAL NOTES 30-Jun-2015 20:02:53 GMT From CASE: 9408548 - Event
The conduits have been proofed and 2 have been confirmed usable. The replacement cable remains en route. The OSP Engineer reported that 2 splice trailers have been requested so that splicing can take place at both ends simultaneously after the replacement cable has been ran. No ETR has been provided.



*** CASCADED EXTERNAL NOTES 30-Jun-2015 20:33:48 GMT From CASE: 9408548 - Event
The Outside Plant Engineer advised that teams are awaiting the arrival of the replacement 88-count fiber cable that is en route from Salem, OR. Splice trailers are on site and placed at each end to begin splicing once the replacement cable is pulled through and prepared for splicing. An estimated arrival time for the replacement cable is unavailable due to heavy traffic in the area. Additionally, an estimated restoral time remains unavailable.


Comment by OVH - Wednesday, 01 July 2015, 13:48PM

*** CASCADED EXTERNAL NOTES 30-Jun-2015 21:41:20 GMT From CASE: 9408548 - Event
The replacement fiber has arrived on site and crews are working to pull the fiber through. Teams estimate that it will take approximately one hour to pull the cable. Once pulled through, splice crews are in place to begin splicing at both ends simultaneously. An estimated restoral time remains unavailable.


Comment by OVH - Wednesday, 01 July 2015, 13:49PM

*** CASCADED EXTERNAL NOTES 30-Jun-2015 22:55:36 GMT From CASE: 9408548 - Event
Field Services advised that the 88-count cable has been pulled through completely and splicers are on both sides working diligently to finish preparing the fiber for splicing. Teams estimate splicing will begin on the 88-count cable in approximately 20 minutes. The 102-count cable is being pulled through at this time and splicing is expected to begin once the 88-count cable has been spliced (approximately 45 minutes).

*** CASCADED EXTERNAL NOTES 30-Jun-2015 23:39:39 GMT From CASE: 9408548 - Event
Splicing has commenced on the 88-count cable but an estimated restoral time for services has not been provided.


Comment by OVH - Wednesday, 01 July 2015, 14:04PM

The router is Up!